PQR Flour Mills Co. is an Omani Company made it’s beginning in the year 1977 with a capacity to mill 150 metric tons per day of wheat flour and allied products. The company gradually increased our flour milling capacity over the years and today has a capacity of 800 Metric Ton per day. In 1983 the company set up an animal feed mill having milling capacity of 300 Metric ton per day. Today the company has a capacity of 1500 Metric ton per day for Animal Feed. Their trading results for the year end
1
Expert's answer
2021-09-17T09:21:04-0400

Solution:

Opening capacity wheat flour = 150 metric tons

Closing capacity wheat flour = 800 metric tons

Profit levels = 800 – 150 = 650

Opening capacity Animal feed = 300 metric tons

Closing capacity Animal feed = 1500 metric tons

Profit levels = 1500 – 300 = 1,200

Total trading results = 650 + 120 = 1,850